### v1

@@ Abstract
Large language models (LLMs), with their powerful generative capabilities and vast knowledge, empower various tasks in everyday life. However, these abilities are primarily concentrated in high-resource languages, leaving low-resource languages with weaker generative capabilities and relatively limited knowledge. Enhancing the multilingual capabilities of LLMs is therefore crucial for serving over 100 linguistic communities worldwide. An intuitive approach to enhance the multilingual capabilities would be to construct instruction data for various languages, but constructing instruction data for over 100 languages is prohibitively costly. In this paper, we introduce BayLing 2, which efficiently transfers generative capabilities and knowledge from high-resource languages to low-resource languages through language alignment. To achieve this, we constructed a dataset of 3.2 million instructions, comprising high-resource language instructions (Chinese and English) and cross-lingual instructions for 100+ languages and performed instruction tuning based on the dataset to facilitate the capability transfer between languages. Using Llama as the foundation model, we developed BayLing-2-7B, BayLing-2-13B, and BayLing-3-8B, and conducted a comprehensive evaluation of BayLing. For multilingual translation across 100+ languages, BayLing shows superior performance compared to open-source models of similar scale. For multilingual knowledge and understanding benchmarks, BayLing achieves significant improvements across over 20 low-resource languages, demonstrating its capability of effective knowledge transfer from high-resource to low-resource languages. Furthermore, results on English benchmarks indicate that BayLing maintains high performance in high-resource languages while enhancing the performance in low-resource languages. Demo [^1] Demo: http://nlp.ict.ac.cn/bayling/demo  3Homepage: http://nlp.ict.ac.cn/bayling, homepage22footnotemark: 2, code [^3] Code: https://github.com/ictnlp/BayLing   5Models: BayLing-2-7B, BayLing-2-13B, BayLing-3-8B and models44footnotemark: 4of BayLing are available.
@@ 1 Introduction
Based on foundational models Llama-2-7B-Chat, Llama-2-13B-Chat, and Llama-3-8B-Instruct, we developed BayLing-2-7B, BayLing-2-13B, and BayLing-3-8B through the proposed efficient language alignment. We conducted a comprehensive evaluation of BayLing’s performance on both multilingual and general tasks and assessed the quality of language alignment through multilingual translation using the Flores-101 and WMT22 benchmarks. BayLing showed superior translation performance across more than 100 languages, achieving the best results among open-source models of comparable scale. We further evaluated BayLing’s multilingual knowledge and generative capabilities using benchmarks including Belebele, Multilingual HellaSwag, XNLI, and Multilingual ARC. The results indicated significant performance improvements across more than 20 low-resource languages, such as Bambara, Luganda, Swahili, and Zulu. This demonstrates effective knowledge and generative capability transfer from high-resource to low-resource languages. Additionally, we evaluated BayLing on various general benchmarks (primarily in English), finding that the language alignment had minimal impacts on BayLing’s performance in high-resource languages.
@@ 3.2 Training
Figure 5: Training loss curve of BayLing-3-8B.
Using Llama-2-7B-Chat, Llama-2-13B-Chat and Llama-3-8B-Instruct as foundational models, We fine-tune BayLing-2-7B, BayLing-2-13B and BayLing-3-8B respectively on the instruction dataset proposed in Section 3.1. We fine-tune BayLing 2 on 8 NVIDIA A800 80G GPUs for 3 epochs, using a global batch size of 128, learning rate of 2e-5 and weight decay of 0.0. Note that we apply learning rate of 2e-6 for BayLing-3-8B. We employ DeepSpeed (Rasley et al., 2020) and Gradient Checkpointing (Chen et al., 2016) techniques to optimize memory consumption. The training loss curve of BayLing-3-8B is depicted in Figure 5.
@@ 4 Evaluation
In this section, we comprehensively evaluate the performance of BayLing-2-7B, BayLing-2-13B and BayLing-3-8B on multilingual tasks and general tasks respectively.
@@ 4.1.2 Multilingual Multi-task Evaluation
Figure 8(a), 8(b), 8(c), 8(d) provide detailed illustrations of the experimental outcomes on the Belebele, Multilingual HellaSwag, XNLI, Multilingual ARC benchmarks across several low-resource languages. The BayLing-2-7B and BayLing-3-8B models demonstrate notable performance benefits. Among these, BayLing-3-8B consistently delivers the best results across most of the low-resource languages evaluated. Meanwhile, BayLing-2-7B outperforms other 7B models in most of these languages. Remarkably, BayLing-2-7B even surpasses the Llama-3-8B-Instruct model in the Swati language subset on Belebele benchmark (belebele_ssw_Latn).
@@ 4.2 General Capability
Figure 9 illustrates the performance of BayLing on the specified benchmarks. BayLing-2-7B and BayLing-3-8B demonstrate exceptional performance across several benchmarks. Notably, BayLing-3-8B outperforms all other models, achieving a score of 0.8214 on the CommitmentBank Benchmark. BayLing-2-7B attains the highest performance on GLUE and GSM8K benchmarks. Despite not being specifically trained for it, BayLing-2-7B and BayLing-3-8B still deliver comparable performances on other benchmarks when compared to other models. Overall, BayLing enhances the multilingual capabilities of LLMs, especially in low-resource languages, without significantly impacting the performance in high-resource languages. This indicates that BayLing effectively mitigates multilingual conflicts within LLM through language alignment.
@@ 4.3 Effect of Language Alignment
To validate the effect of language alignment brought by cross-lingual instructions, we conducted an ablation study on cross-lingual instructions. Specifically, we removed all cross-lingual instructions from the training data, denoting this variant as BayLing-3-8B (w/o cross-lingual instructions).
Improving Performance of Low-resource Languages Figure 10 compares the performance of BayLing-3-8B and BayLing-3-8B (w/o cross-lingual instructions) on the multilingual benchmark Belebele. The results show that cross-lingual instructions significantly enhance LLM performance in low-resource languages. This indicates that cross-lingual instructions successfully help LLM achieve language alignment, thereby transferring knowledge and comprehension from high-resource languages to low-resource ones. When removing all cross-lingual instructions, the performance of LLMs in low-resource languages is adversely affected due to catastrophic forgetting. Therefore, involving cross-lingual instructions in supervised fine-tuning is both efficient and crucial for improving the multilingual capabilities of LLMs.
Avoiding Inter-language Conflicts Figure 11 compares BayLing-3-8B and BayLing-3-8B (w/o cross-lingual instructions) on the Chinese/English benchmark. When removing cross-lingual instructions, we observed a significant performance decline in Chinese benchmark, indicating LLM will suffer from conflicts between Chinese and English instructions. The presence of numerous cross-lingual instructions between Chinese and English largely prevents these conflicts. Therefore, to simultaneously enhance LLM performance across multiple languages, introducing cross-lingual instructions is an effective way to avoid inter-language conflicts.
@@ Acknowledgements
We extend our heartfelt gratitude to everyone who contributed to the development of BayLing 2. In particular, we would like to thank Ms. Xiaohong Wang for her insightful feedback and valuable suggestions regarding the use of InforSuperBahn MLOps, as well as for her exceptional support in organizing resources, providing computational infrastructure, and facilitating the presentation of BayLing 2. We also acknowledge the dedicated efforts of the development team at the Nanjing Institute of InforSuperBahn for their indispensable role in maintaining computational resources and designing the display interface for BayLing 2’s webpage and demonstration.
@@ Appendix A Flores-101 Benchmark
Table 3, 4, 5, and 6 report the numerical results of Llama-3-8B-Instruct and BayLing-3-8B on the Flores-101 benchmark.
@@
Table 4: BLEU scores of BayLing-3-8B on Flores-101 benchmark.
BayLing-3-8B
@@
Table 6: COMET scores of BayLing-3-8B on Flores-101 benchmark.
BayLing-3-8B

### v3

@@ Abstract
Large language models (LLMs), with their powerful generative capabilities and vast knowledge, empower various tasks in everyday life. However, these abilities are primarily concentrated in high-resource languages, leaving low-resource languages with weaker generative capabilities and relatively limited knowledge. Enhancing the multilingual capabilities of LLMs is therefore crucial for serving over 100 linguistic communities worldwide. An intuitive approach to enhance the multilingual capabilities would be to construct instruction data for various languages, but constructing instruction data for over 100 languages is prohibitively costly. In this paper, we introduce BayLing 2, which efficiently transfers generative capabilities and knowledge from high-resource languages to low-resource languages through language alignment. To achieve this, we constructed a dataset of 3.2 million instructions, comprising high-resource language instructions (Chinese and English) and cross-lingual instructions for 100+ languages and performed instruction tuning based on the dataset to facilitate the capability transfer between languages. Using Llama as the foundation model, we developed BayLing-2-7B, BayLing-2-13B, and BayLing-2-8B, and conducted a comprehensive evaluation of BayLing. For multilingual translation across 100+ languages, BayLing shows superior performance compared to open-source models of similar scale. For multilingual knowledge and understanding benchmarks, BayLing achieves significant improvements across over 20 low-resource languages, demonstrating its capability of effective knowledge transfer from high-resource to low-resource languages. Furthermore, results on English benchmarks indicate that BayLing maintains high performance in high-resource languages while enhancing the performance in low-resource languages. Demo [^1] Demo: http://nlp.ict.ac.cn/bayling/demo  3Homepage: http://nlp.ict.ac.cn/bayling, homepage22footnotemark: 2, code [^3] Code: https://github.com/ictnlp/BayLing   5Models: BayLing-2-7B, BayLing-2-13B, BayLing-2-8B and models44footnotemark: 4of BayLing are available.
@@ 1 Introduction
Based on foundational models Llama-2-7B-Chat, Llama-2-13B-Chat, and Llama-3-8B-Instruct, we developed BayLing-2-7B, BayLing-2-13B, and BayLing-2-8B through the proposed efficient language alignment. We conducted a comprehensive evaluation of BayLing’s performance on both multilingual and general tasks and assessed the quality of language alignment through multilingual translation using the Flores-101 and WMT22 benchmarks. BayLing showed superior translation performance across more than 100 languages, achieving the best results among open-source models of comparable scale. We further evaluated BayLing’s multilingual knowledge and generative capabilities using benchmarks including Belebele, Multilingual HellaSwag, XNLI, and Multilingual ARC. The results indicated significant performance improvements across more than 20 low-resource languages, such as Bambara, Luganda, Swahili, and Zulu. This demonstrates effective knowledge and generative capability transfer from high-resource to low-resource languages. Additionally, we evaluated BayLing on various general benchmarks (primarily in English), finding that the language alignment had minimal impacts on BayLing’s performance in high-resource languages.
@@ 3.2 Training
Figure 5: Training loss curve of BayLing-2-8B.
Using Llama-2-7B-Chat, Llama-2-13B-Chat and Llama-3-8B-Instruct as foundational models, We fine-tune BayLing-2-7B, BayLing-2-13B and BayLing-2-8B respectively on the instruction dataset proposed in Section 3.1. We fine-tune BayLing 2 on 8 NVIDIA A800 80G GPUs for 3 epochs, using a global batch size of 128, learning rate of 2e-5 and weight decay of 0.0. Note that we apply learning rate of 2e-6 for BayLing-2-8B. We employ DeepSpeed (Rasley et al., 2020) and Gradient Checkpointing (Chen et al., 2016) techniques to optimize memory consumption. The training loss curve of BayLing-2-8B is depicted in Figure 5.
@@ 4 Evaluation
In this section, we comprehensively evaluate the performance of BayLing-2-7B, BayLing-2-13B and BayLing-2-8B on multilingual tasks and general tasks respectively.
@@ 4.1.2 Multilingual Multi-task Evaluation
Figure 8(a), 8(b), 8(c), 8(d) provide detailed illustrations of the experimental outcomes on the Belebele, Multilingual HellaSwag, XNLI, Multilingual ARC benchmarks across several low-resource languages. The BayLing-2-7B and BayLing-2-8B models demonstrate notable performance benefits. Among these, BayLing-2-8B consistently delivers the best results across most of the low-resource languages evaluated. Meanwhile, BayLing-2-7B outperforms other 7B models in most of these languages. Remarkably, BayLing-2-7B even surpasses the Llama-3-8B-Instruct model in the Swati language subset on Belebele benchmark (belebele_ssw_Latn).
@@ 4.2 General Capability
Figure 9 illustrates the performance of BayLing on the specified benchmarks. BayLing-2-7B and BayLing-2-8B demonstrate exceptional performance across several benchmarks. Notably, BayLing-2-8B outperforms all other models, achieving a score of 0.8214 on the CommitmentBank Benchmark. BayLing-2-7B attains the highest performance on GLUE and GSM8K benchmarks. Despite not being specifically trained for it, BayLing-2-7B and BayLing-2-8B still deliver comparable performances on other benchmarks when compared to other models. Overall, BayLing enhances the multilingual capabilities of LLMs, especially in low-resource languages, without significantly impacting the performance in high-resource languages. This indicates that BayLing effectively mitigates multilingual conflicts within LLM through language alignment.
@@ 4.3 Effect of Language Alignment
To validate the effect of language alignment brought by cross-lingual instructions, we conducted an ablation study on cross-lingual instructions. Specifically, we removed all cross-lingual instructions from the training data, denoting this variant as BayLing-2-8B (w/o cross-lingual instructions).
Improving Performance of Low-resource Languages Figure 10 compares the performance of BayLing-2-8B and BayLing-2-8B (w/o cross-lingual instructions) on the multilingual benchmark Belebele. The results show that cross-lingual instructions significantly enhance LLM performance in low-resource languages. This indicates that cross-lingual instructions successfully help LLM achieve language alignment, thereby transferring knowledge and comprehension from high-resource languages to low-resource ones. When removing all cross-lingual instructions, the performance of LLMs in low-resource languages is adversely affected due to catastrophic forgetting. Therefore, involving cross-lingual instructions in supervised fine-tuning is both efficient and crucial for improving the multilingual capabilities of LLMs.
Avoiding Inter-language Conflicts Figure 11 compares BayLing-2-8B and BayLing-2-8B (w/o cross-lingual instructions) on the Chinese/English benchmark. When removing cross-lingual instructions, we observed a significant performance decline in Chinese benchmark, indicating LLM will suffer from conflicts between Chinese and English instructions. The presence of numerous cross-lingual instructions between Chinese and English largely prevents these conflicts. Therefore, to simultaneously enhance LLM performance across multiple languages, introducing cross-lingual instructions is an effective way to avoid inter-language conflicts.
@@ Acknowledgements
We extend our heartfelt gratitude to everyone who contributed to the development of BayLing 2. In particular, we would like to thank Ms. Xiaohong Wang for her insightful feedback and valuable suggestions regarding the use of OneAiNexus, as well as for her exceptional support in organizing resources, providing computational infrastructure, and facilitating the presentation of BayLing 2. Besides, we would like to express thanks to the Sothis.AI for the support in training of BayLing 2. Special thanks are due to the team of Nanjing Institute of InforSuperBahn - Intelligent Computility Platform Research Center, who played an indispensable role in maintaining the computational resources, designing the BayLing 2’s webpage and demonstrating the system interface.
@@ Appendix A Flores-101 Benchmark
Table 3, 4, 5, and 6 report the numerical results of Llama-3-8B-Instruct and BayLing-2-8B on the Flores-101 benchmark.
@@
Table 4: BLEU scores of BayLing-2-8B on Flores-101 benchmark.
BayLing-2-8B
@@
Table 6: COMET scores of BayLing-2-8B on Flores-101 benchmark.
BayLing-2-8B